# Disentanglement in Implicit Causal Models via Switch Variable

### v1

@@ 4 Experiments and Results
The experiments conducted in this paper address two downstream tasks; (1) Causal Disentanglement to identify the true causal graph from pairs of observations $(x,\tilde{x})$, and (2) Action Inference to make supervised inferences about actions generated from the post-intervention samples using information about the values of the manipulated causal variables.
@@ 4.1.1 Synthetic Dataset
4.1.2 Action Datasets
Causal-Triplet datasets are designed specifically for actionable counterfactuals Liu et al. (2023) with paired images where few or more global scene properties might change such as camera view and object occlusions. Therefore, the images can be seen as a result of soft intervention where an action is performed on an object but other subtle changes are also allowed. These datasets Liu et al. (2023) consist of: images obtained from a photo-realistic simulator of embodied agents, ProcTHOR Deitke et al. (2022), and the other contains images repurposed from a real-world video dataset of human-object interactions Damen et al. (2022). The former one contains 100k images in which 7 types of actions manipulate 24 types of objects in 10k distinct ProcTHOR indoor environments. The latter consists of 2632 image pairs, collected under a similar setup from the Epic-Kitchens dataset with 97 actions manipulating 277 objects.
@@ 4.2 Metrics
For the causal disentanglement task, we are going to use the DCI scores Eastwood and Williams (2018). Causal disentanglement score quantifies the degree to which $\mathcal{Z}_{i}$ factorises or disentangles the $\mathcal{Z}^{*}$. Causal disentanglement $D_{i}$ for $\mathcal{Z}_{i}$ is calculated as $D_{i}=(1-H_{K}(P_{i.}))=(1+\sum_{k=0}^{K-1}P_{ik}\log_{K}P_{ik})$ where $P_{ij}=\frac{R_{ij}}{\sum_{k=0}^{K-1}R_{ik}}$ and $R_{ij}$ denotes the probability of $\mathcal{Z}_{i}$ being important for predicting $\mathcal{Z}_{j}^{*}$. Total causal disentanglement is the weighted average $\sum_{i}\rho_{i}D_{i}$ where $\rho_{i}=\frac{\sum_{j}R_{ij}}{\sum_{ij}R_{ij}}$. Causal Completeness quantifies the degree to which each $\mathcal{Z}_{i}^{*}$ is captured by a single $\mathcal{Z}_{i}$. Causal completeness is calculated as $C_{j}=(1-H_{D}(\tilde{P}_{.j}))=(1+\sum_{d=0}^{D-1}\tilde{P}_{dj}\log_{D}\tilde{P}_{ij})$. $D$ and $K$ here are equal to the dimension of $\mathcal{Z}^{*}$ and $\mathcal{Z}$ which is $n$. For the action inference task, we will use classification accuracy as a metric.
@@ 5.1 Causal Disentanglement
We generated a dataset for the soft interventions and trained the models of SoftCD, ILCM, $\beta$-VAE and D-VAE for 10 different seeds, which generated 10 different causal graphs. We selected 4 causal variables to encompass complex causal structures, including forks, chains, and colliders. Table 2 displays the Causal Disentanglement and Causal Completeness scores for all models, computed on the test data.
@@
Graph Causal Disentanglement Causal Completeness Model Name $\beta$-VAE $d$-VAE ILCM SoftCD $\beta$-VAE $d$-VAE ILCM SoftCD G1 0.39 0.54 0.69 0.81 0.53 0.68 0.76 0.87 G2 0.15 0.72 0.74 0.83 0.24 0.77 0.79 0.87 G3 0.25 0.51 0.67 0.97 0.54 0.56 0.77 0.97 G4 0.19 0.50 0.65 0.69 0.41 0.69 0.76 0.79 G5 0.25 0.44 0.32 0.42 0.45 0.54 0.37 0.50 G6 0.51 0.62 0.73 0.98 0.63 0.69 0.73 0.98 G7 0.35 0.49 0.70 0.76 0.65 0.74 0.88 0.90 G8 0.48 0.54 0.50 0.60 0.61 0.63 0.62 0.69 G9 0.31 0.68 0.83 0.86 0.40 0.76 0.86 0.87 G10 0.15 0.39 0.52 0.32 0.42 0.56 0.82 0.50
The results in Table 2 indicate that our method SoftCD can identify the true causal graph in most cases. The worst results are seen for graphs $G5$ and $G10$. As mentioned in Schölkopf et al. (2021); Perry et al. (2022), causal graphs are sparse and in the $G5$ case, where the graph is fully connected, the proposed method cannot identify the causal variables well. Furthermore, in the next experiment we are going to examine the factors affecting causal disentanglement such as the number of edges in the graph and the intensity of soft intervention effect. These findings can explain why SoftCD cannot identify causal variables in $G10$ despite its sparsity.
Table 3: Table comparing action and object accuracy across various methods on Causal-Triplet datasets under different settings. $Z$ and $z_{i}$ show whether all causal variables ($Z$), or only the intervened casual variable ($z_{i}$) are used for the prediction task. $R_{64}$ and $R_{32}$ denote images with resolutions $64\times 64$ and $32\times 32$, respectively.
@@
Method $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $z_{i}$;$R_{32}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $z_{i}$;$R_{32}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $z_{i}$;$R_{32}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $z_{i}$;$R_{32}$
$\beta-VAE$ Higgins et al. (2017) 0.30 0.17 0.13 0.20 0.05 0.05 0.32 0.20 0.22 0.41 0.35 0.35
$d-VAE$ Locatello et al. (2020) 0.19 0.60 0.46 0.20 0.14 0.1 0.27 0.60 0.63 0.38 0.77 0.78
ILCM Brehmer et al. (2022) 0.17 0.25 0.27 0.18 0.06 0.07 0.21 0.36 0.35 0.34 0.54 0.59
SoftCD (ours) 0.17 0.63 0.66 0.21 0.14 0.16 0.22 0.75 0.95 0.36 0.78 0.81
@@ 5.2 Factors Affecting Causal Disentanglement
Table 4: SoftCD performance on different configurations of $G5$
Table 5: Action and object accuracy of two explicit models are compared with SoftCD. Experiments are conducted applying image with resolution of $R_{64}$ as the input to the encoder with the intervened casual variable ($z_{i}$).
Datasets Methods Action Accuracy Object Accuracy
Epic-Kitchens ENCO Lippe et al. (2022b) 0.39 0.07
DDS Charpentier et al. (2022) 0.49 0.09
Fixed-order 0.34 0.09
SoftCD (ours) 0.63 0.14
ProcTHOR ENCO Lippe et al. (2022b) 0.60 0.64
DDS Charpentier et al. (2022) 0.38 0.44
Fixed-order 0.43 0.55
SoftCD (ours) 0.75 0.78
The results in Table 4 further confirms the sparsity of causal graphs as the causal disentanglement is much worse in the fully-connected graph than the default graph of $G3$. The result for significantly different post-intervention causal mechanisms indicate that the switch variable cannot approximate intense effects of soft intervention and more supervision is required to observe $\mathcal{V}$. Similar post-intervention causal mechanisms also do not have sufficient variability to disentangle the causal variables as mentioned in Theory 3.9.
@@ 5.3 Action Inference
In this experiment, we show the performance of SoftCD in the real-world Causal-Triplet datasets. Based on the nature of actions in this dataset, the causal variables should represent attributes of objects such as shape and color. As the dataset consists of images we train all the methods with ResNet encoder and decoder. One of the constraints of ILCM and SoftCD is to have a diffeomorphic encoder.
To ensure this, we add the following inverse consistency constraint to the loss function to guarantee that decoder is the inverse of encoder as: $\sum_{i}E_{e_{i}\sim q(e_{i}|x),\hat{e}_{i}\sim q(\hat{e}_{i}|\hat{x})}[(e_{i}-\hat{e}_{i})^{2}]$, where $\hat{x}$ are the reconstructed samples. For the ProcThor dataset the number of causal variables are 7. For the Epic-Kitchens dataset, we randomly chose 20 actions from the dataset as 97 causal variables will be too complex in a VAE setup. The results are shown in Table 3.
The results in Table 3 indicate that when including all causal variables to predict actions, SoftCD performs at par with the baseline methods. However, including all causal variables in the action or object inference may cause spurious correlations. Therefore, we have also experimented with including only the related causal variable in action and object inference. In this setting, SoftCD significantly outperforms the baseline methods which means that it can better disentangle the causal variables.
We have also compared SoftCD with explicit causal representation learning methods. ENCO Lippe et al. (2022b) and DDS Charpentier et al. (2022) have variable topological order of causal variables during training. Furthermore, we have included a specific setting where the topological order is fixed during training. As shown in Table 5, our proposed method has superior performance to explicit models as well.
@@ 6 Conclusion
We have presented SoftCD, a novel model, which improves implicit causal representation learning during soft interventions by introducing a causal mechanism switch variable. In evaluations on synthetic and real-world datasets, SoftCD outperforms state-of-the-art methods, demonstrating its effectiveness in practical applications. Theoretical contributions include articulating conditions for achieving causal disentanglement in implicit models through soft interventions. Our results underscore SoftCD’s capacity to discern causal models from soft interventions, establishing it as a promising avenue for future research.
Appendix A1 Generative model
Figure A1 shows our main generative model. It includes a data augmentation step that adds the intervention displacement $\tilde{x}-x$ as an observed feature that directly represents the effect of an intervention in observation space.
Figure A1: Our generative model. $x,\tilde{x}$ represent low-level observed variables. $e,\tilde{e}$ represent higher-level latent exogenous variables, and $z,\tilde{z}$ represent higher-level latent causal variables (e.g. objects). $v$ represents the causal mechanism switch variable.
@@
(Diffeomorphism) A diffeomorphism between smooth manifolds $M$ and $N$ is a bijective map $f:M\rightarrow N$, which is smooth and has a smooth inverse.
Diffeomorphisms preserve information as they are invertible transformations without discontinuous changes in their image.
@@
In other words, the pushforward measure $f_{*}p$ assigns a measure to a set in $B$ by measuring the pre-image of that set under $f$ in the space $A$.
@@
(Identifiability of augmented latent causal models) Let $M=(A,X,g,I)$ and $M^{\prime}=(A^{\prime},X^{\prime},g^{\prime},I^{\prime})$ be two Augmented Latent Causal Models (ALCM) with identical observation space $X=X^{\prime}$ and identical intervention targets $I=I^{\prime}$.
If the following conditions are satisfied,
  • Assumptions for identical correspondence as explained in Assumption 3.4.
  • Some form of supervision on how the soft interventions affect the causal mechanisms or observability of $V$ as explained in Assumption 3.8.
  • The causal and exogenous variables are real-valued $Z_{i}=Z^{\prime}_{i}=E_{i}=E^{\prime}_{i}=\mathbb{R}$.
  • The causal mechanism switch variables are real valued $V=\mathbb{R}^{n}$.
@@
  1. 1.
    The ALCMs assign the same likelihood to interventional and observational data i.e., $p^{X}_{M}(x,\tilde{x})=p^{X}_{M^{\prime}}(x,\tilde{x})$.
  2. 2.
    $M\sim_{r}M^{\prime}$ according to Definition 3.6.

### v2

@@ 4 Experiments and Results
The experiments conducted in this paper address two downstream tasks; (1) Causal Disentanglement to identify the true causal graph from pairs of observations $(x,\tilde{x},i)$, and (2) Action Inference to make supervised inferences about actions generated from the post-intervention samples using information about the values of the manipulated causal variables. Moreover, we conducted additional experiments designed as an ablation study, the results of which are presented in A4. All models are trained using the same setting and data with known intervention targets.
@@ 4.1 Datasets
Action Datasets
Causal-Triplet datasets tailored for actionable counterfactuals [19] feature paired images where several global scene properties may vary including camera view and object occlusions. Thus, the images can be viewed as outcomes of soft interventions, wherein actions affect objects alongside subtle alterations. These datasets [19] consist of: images obtained from a photo-realistic simulator of embodied agents, ProcTHOR [9], and the other contains images repurposed from a real-world video dataset of human-object interactions [8]. The former one contains 100 k images in which 7 types of actions manipulate 24 types of objects in 10 k distinct ProcTHOR indoor environments. The latter consists of 2,632 image pairs, collected under a similar setup from the Epic-Kitchens dataset with 97 actions manipulating 277 objects.Based on the nature of actions in this dataset, the causal variables should represent attributes of objects such as shape and color. As the dataset consists of images we train all the methods with ResNet encoder and decoder. For the ProcThor dataset the number of causal variables are 7. For the Epic-Kitchens dataset, we randomly chose 20 actions from the dataset as 97 causal variables will be too complex in a VAE setup.
@@ 4.2 Metrics
For the causal disentanglement task, we are going to use the DCI scores [10]. Causal disentanglement score quantifies the degree to which $\mathcal{Z}_{i}$ factorises or disentangles the $\mathcal{Z}^{*}$. Causal disentanglement $D_{i}$ for $\mathcal{Z}_{i}$ is calculated as $D_{i}=(1-H_{K}(P_{i.}))=(1+\sum_{k=0}^{K-1}P_{ik}\log_{K}P_{ik})$ where $P_{ij}=\frac{R_{ij}}{\sum_{k=0}^{K-1}R_{ik}}$ and $R_{ij}$ denotes the probability of $\mathcal{Z}_{i}$ being important for predicting $\mathcal{Z}_{j}^{*}$. Total causal disentanglement is the weighted average $\sum_{i}\rho_{i}D_{i}$ where $\rho_{i}=\frac{\sum_{j}R_{ij}}{\sum_{ij}R_{ij}}$. Causal Completeness quantifies the degree to which each $\mathcal{Z}_{i}^{*}$ is captured by a single $\mathcal{Z}_{i}$. Causal completeness is calculated as $C_{j}=(1-H_{D}(\tilde{P}_{.j}))=(1+\sum_{d=0}^{D-1}\tilde{P}_{dj}\log_{D}\tilde{P}_{ij})$. $D$ and $K$ here are equal to the dimension of $\mathcal{Z}^{*}$ and $\mathcal{Z}$ which is $n$. For the action inference task, we will use classification accuracy as a metric. As we assume intervention targets are known, we train all models using known intervention targets for a fair comparison.
@@ 5.1 Causal Disentanglement
We generated a dataset for the soft interventions and trained the models of ICRL-SM, ILCM, $\beta$-VAE and D-VAE for 10 different seeds, which generated 10 different causal graphs. We selected 4 causal variables to encompass complex causal structures, including forks, chains, and colliders. Table 2 displays the Causal Disentanglement and Causal Completeness scores for all models, computed on the test data.
Graph Causal Disentanglement Causal Completeness Model Name $\beta$-VAE $d$-VAE ILCM ICRL-SM $\beta$-VAE $d$-VAE ILCM ICRL-SM G1 0.38 0.54 0.71 0.82 0.51 0.69 0.78 0.87 G2 0.30 0.72 0.75 0.83 0.49 0.77 0.80 0.87 G3 0.28 0.51 0.68 0.98 0.49 0.56 0.78 0.98 G4 0.16 0.50 0.65 0.68 0.38 0.69 0.77 0.78 G5 0.27 0.44 0.53 0.42 0.45 0.54 0.66 0.50 G6 0.52 0.62 0.71 0.98 0.66 0.69 0.86 0.98 G7 0.39 0.49 0.71 0.75 0.70 0.73 0.89 0.89 G8 0.47 0.54 0.50 0.59 0.6 0.63 0.62 0.68 G9 0.30 0.68 0.83 0.85 0.40 0.76 0.86 0.87 G10 0.39 0.39 0.52 0.32 0.53 0.56 0.82 0.70
@@
The results in Table 2 indicate that our method ICRL-SM can identify the true causal graph in most cases. The worst results are seen for graphs $G5$ and $G10$. As mentioned in [27, 25], causal graphs are sparse and in the $G5$ case, where the graph is fully connected, the proposed method cannot identify the causal variables well. Furthermore, in the next experiment we are going to examine the factors affecting causal disentanglement such as the number of edges in the graph and the intensity of soft intervention effect. These findings can explain why ICRL-SM cannot identify causal variables in $G10$ despite its sparsity.
Table 3: Table comparing action and object accuracy across various methods on Causal-Triplet datasets under different settings. $Z$ and $z_{i}$ show whether all causal variables ($Z$), or only the intervened casual variable ($z_{i}$) are used for the prediction task. $R_{64}$ denote images with resolutions $64\times 64$.
@@
Method $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$ $Z$;$R_{64}$ $z_{i}$;$R_{64}$
$\beta-VAE$ [11] 0.27 0.18 0.19 0.06 0.39 0.30 0.44 0.37
$d-VAE$ [21] 0.19 0.69 0.20 0.17 0.35 0.81 0.40 0.78
ILCM [3] 0.21 0.59 0.14 0.14 0.30 0.70 0.41 0.76
ICRL-SM (ours) 0.16 0.86 0.16 0.18 0.28 0.93 0.40 0.82
@@ 5.2 Factors Affecting Causal Disentanglement
Action and object accuracy
Datasets Methods Action Accuracy Object Accuracy
Epic-Kitchens ENCO [16] 0.69 0.13
DDS [5] 0.44 0.09
Fixed-order 0.79 0.14
ICRL-SM (ours) 0.86 0.18
ProcTHOR ENCO [16] 0.45 0.53
DDS [5] 0.64 0.67
Fixed-order 0.65 0.54
ICRL-SM (ours) 0.93 0.82
@@
ICRL-SM performance on different configurations of $G5$
Table 4: Left table depicts the action and object accuracy of three explicit models, with experiments conducted applying an image with resolution of $R_{64}$ as the input to the Resnet50 encoder with the intervened causal variable ($z_{i}$). Right table shows the comparison of ICRL-SM performance on different configurations of $G5$
The results in Table 4 further confirms the sparsity of causal graphs as the causal disentanglement is much worse in the fully-connected graph than the default graph of $G3$. The result for significantly different post-intervention causal mechanisms indicate that the switch variable cannot approximate intense effects of soft intervention and more supervision is required to observe $\mathcal{V}$. Similar post-intervention causal mechanisms also do not have sufficient variability to disentangle the causal variables as mentioned in Theory 3.5.
@@ 5.3 Action Inference
In this experiment, we show the performance of ICRL-SM in the real-world Causal-Triplet datasets. In these datasets $\mathcal{V}$ i.e., soft intervention effects, are not directly observable. Nevertheless, our findings suggest that incorporating soft interventions through $\mathcal{V}$ leads to superior performance compared to other implicit modeling approaches. Clearly, understanding the impact of soft interventions on the generative system of the dataset would result in improved outcomes.
The results in Table 3 indicate that when including all causal variables to predict actions, ICRL-SM performs at par with the baseline methods. However, including all causal variables in the action or object inference may cause spurious correlations. Therefore, we have also experimented with including only the related causal variable in action and object inference. In this setting, ICRL-SM significantly outperforms the baseline methods which means that it can better disentangle the causal variables. We have also compared ICRL-SM with explicit causal representation learning methods. ENCO [16] and DDS [5] have variable topological order of causal variables during training. Furthermore, we have included a specific setting where the topological order is fixed during training. As shown in Table 4, our proposed method has superior performance to explicit models as well.
@@ 6 Conclusion
ICRL-SM, our novel model, enhances implicit causal representation learning during soft interventions by introducing a causal mechanism switch variable. Evaluations on synthetic and real-world datasets demonstrate ICRL-SM’s superiority over state-of-the-art methods, highlighting its practical effectiveness. Our findings emphasize ICRL-SM’s ability to discern causal models from soft interventions, marking it as a promising avenue for future research.
@@
Structural Causal Models
A structural causal model (SCM) is a tuple $\mathcal{C}=(\mathcal{F},\mathcal{Z},\mathcal{E},\mathcal{G})$ with the following components:
1. The domain of causal variables $\mathcal{Z}=\mathcal{Z}_{1}\times\mathcal{Z}_{2}\times\ldots\times\mathcal{Z}_{n}$.
2. The domain of exogenous variables $\mathcal{E}=\mathcal{E}_{1}\times\mathcal{E}_{2}\times\ldots\times\mathcal{E}_{n}$.
3. A directed acyclic graph $\mathcal{G}(\mathcal{C})$ over the causal and exogenous variables.
4. A causal mechanism $f_{i}\in\mathcal{F}$ which maps an assignment of parent values for the parents $\mathcal{Z}_{pa_{i}}$ plus an exogenous variable value for $\mathcal{E}_{i}$ to a value of causal variable $Z_{i}$.
Definition A1.2.
(Component-wise Transformation) Let $\phi$ be a transformation (1-1 onto mapping) between product spaces $\phi:\Pi_{i=1}^{n}\mathcal{X}_{i}\rightarrow\Pi_{i=1}^{n}\mathcal{Y}_{i}$. If there exist local transformations $\phi_{i}$ such that $\forall i,j,\>\forall x,\>\phi(x_{1},x_{2},...,x_{n})_{i}=\phi_{i}(x_{j})$, then $\phi$ is a component-wise transformation.
Definition A1.3.
(Diffeomorphism) A diffeomorphism between smooth manifolds $M$ and $N$ is a bijective map $f:M\rightarrow N$, which is smooth and has a smooth inverse. Diffeomorphisms preserve information as they are invertible transformations without discontinuous changes in their image.
@@
where denotes the pushforward operation. In other words, the pushforward measure $f_{*}p$ assigns a measure to a set in $B$ by measuring the pre-image of that set under $f$ in the space $A$.
Lemma A1.5.
The transformation $\phi_{\mathcal{Z}}:\mathcal{Z}\rightarrow\mathcal{Z^{\prime}}$ between the causal variable of two LCMs $\mathcal{M}$ and $\mathcal{M^{\prime}}$ defined in Definition 3.4 is a component-wise transformation, if $\>\forall i,j,i\neq j\quad\tilde{\mathcal{E}^{\prime}_{i}}\perp\!\!\!\perp\tilde{\mathcal{E}^{\prime}_{j}}$ and the causal variables follow a multivariate normal distribution conditional on the pre-intervention exogenous variables where $\tilde{E^{\prime}_{i}}$ denote the post-intervention exogenous variable of causal variable $i$ in $\mathcal{M^{\prime}}$.
proof: We consider the case where the exogenous variables are mapped to causal variables by a location-scale noise model such that $\tilde{z}_{i}=\frac{\tilde{e}_{i}-\widetilde{loc}(e_{/i})}{\widetilde{scale}(e_{/i})}$.
$\forall i,j,i\neq j\quad\tilde{\mathcal{E}^{\prime}_{i}}\perp\!\!\!\perp\tilde{\mathcal{E}^{\prime}_{j}}\rightarrow E[\tilde{\mathcal{E}^{\prime}_{i}}\tilde{\mathcal{E}^{\prime}_{j}}]=E[\tilde{\mathcal{E}^{\prime}_{i}}]E[\tilde{\mathcal{E}^{\prime}_{j}}]$
let’s add these three constants $-E[\tilde{\mathcal{E}^{\prime}_{i}}]\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j}),\>-E[\tilde{\mathcal{E}^{\prime}_{j}}]\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i}),\>\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})$ to the both sides of the equality and then divide both sides by $\widetilde{scale^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{scale^{\prime}_{j}}(e^{\prime}_{/j})$:
$E\left[\frac{\tilde{\mathcal{E}^{\prime}_{i}}\tilde{\mathcal{E}^{\prime}_{j}}-\tilde{\mathcal{E}^{\prime}_{i}}\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})-\tilde{\mathcal{E}^{\prime}_{j}}\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})+\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})}{\widetilde{scale^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{scale^{\prime}_{j}}(e^{\prime}_{/j})}\right]=$
$\frac{E[\tilde{\mathcal{E}^{\prime}_{i}}]E[\tilde{\mathcal{E}^{\prime}_{j}}]-E[\tilde{\mathcal{E}^{\prime}_{i}}]\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})-E[\tilde{\mathcal{E}^{\prime}_{j}}]\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})+\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})}{\widetilde{scale^{\prime}_{i}}(e^{\prime}_{/i})\widetilde{scale^{\prime}_{j}}(e^{\prime}_{/j})}$
$\rightarrow E\left[(\frac{\tilde{\mathcal{E}^{\prime}_{i}}-\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})}{\widetilde{scale^{\prime}_{i}}(e^{\prime}_{/i})})(\frac{\tilde{\mathcal{E}^{\prime}_{j}}-\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})}{\widetilde{scale^{\prime}_{j}}(e^{\prime}_{/j})})\right]=(\frac{E[\tilde{\mathcal{E}^{\prime}_{i}}]-\widetilde{loc^{\prime}_{i}}(e^{\prime}_{/i})}{\widetilde{scale^{\prime}_{i}}(e^{\prime}_{/i})})(\frac{E[\tilde{\mathcal{E}^{\prime}_{j}}]-\widetilde{loc^{\prime}_{j}}(e^{\prime}_{/j})}{\widetilde{scale^{\prime}_{j}}(e^{\prime}_{/j})})$
$\rightarrow E[\tilde{\mathcal{Z}^{\prime}_{i}}\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]=E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]$
$\rightarrow E[\tilde{\mathcal{Z}^{\prime}_{i}}\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]-E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]=0$
$\rightarrow E[\tilde{\mathcal{Z}^{\prime}_{i}}\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]-E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]-E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]+E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]=0$
$\rightarrow E[\tilde{\mathcal{Z}^{\prime}_{i}}\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]-E[\tilde{\mathcal{Z}^{\prime}_{j}}E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]|\mathcal{E}^{\prime}]-E[\tilde{\mathcal{Z}^{\prime}_{i}}E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]|\mathcal{E}^{\prime}]+E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}]E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}]=0$
$\rightarrow E\left[(\tilde{\mathcal{Z}^{\prime}_{i}}-E[\tilde{\mathcal{Z}^{\prime}_{i}}|\mathcal{E}^{\prime}])(\tilde{\mathcal{Z}^{\prime}_{j}}-E[\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}])|\mathcal{E}^{\prime}\right]=0$
$\rightarrow\text{cov}(\tilde{\mathcal{Z}^{\prime}_{i}},\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime})=0$
Typically, the aforementioned equalities would be valid for any diffeomorphic solution function $\tilde{s}_{i}:\tilde{\mathcal{E}}_{i}\rightarrow\tilde{\mathcal{Z}}_{i}$. However, in this paper, we specifically focus on solution functions represented by a location-scale noise model.
Assuming that the causal variables follow a multivariate normal distribution conditional on the pre-intervention exogenous variables, $\text{cov}(\tilde{Z^{\prime}_{i}},\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime})=0$ would imply that $\tilde{\mathcal{Z}^{\prime}_{i}}\perp\!\!\!\perp\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E}^{\prime}$. Let’s define $\phi_{\mathcal{E}}=g^{\prime-1}\circ g:\mathcal{E}\rightarrow\mathcal{E^{\prime}}$ where $g$ and $g^{\prime}$ are the decoders in $\mathcal{M}$ and $\mathcal{M^{\prime}}$. As stated in Assumption 3.1, the decoders are diffeomorphism, hence, $\phi_{\mathcal{E}}$ is a diffeomorphism. Furthermore, let’s denote $\tilde{s}$ as the set of all solution functions in post-intervention which are also diffeomorphism as stated in Assumption 3.1. Consequently:
[…]
$\rightarrow p(\tilde{\mathcal{Z}_{i}^{\prime}}|\mathcal{E})p(\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E})=p(\tilde{\mathcal{Z}_{i}^{\prime}},\tilde{\mathcal{Z}^{\prime}_{j}}|\mathcal{E})$
(all functions in s~ are $(\text{all functions in $\tilde{s}$ are diffeomorphism})\rightarrow p(\tilde{\mathcal{Z}_{i}^{\prime}}|\tilde{s}(\mathcal{E}))p(\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{s}(E))=p(\tilde{\mathcal{Z}_{i}^{\prime}},\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{s}(\mathcal{E}))$
$\rightarrow p(\tilde{\mathcal{Z}_{i}^{\prime}}|\tilde{\mathcal{Z}})p(\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{\mathcal{Z}})=p(\tilde{\mathcal{Z}_{i}^{\prime}},\tilde{\mathcal{Z}^{\prime}_{j}}|\tilde{\mathcal{Z}})$
The association between $\tilde{\mathcal{Z}}^{\prime}$ and $\tilde{\mathcal{Z}}$ arises from their shared observation space. We know that every causal variable in $\mathcal{M^{\prime}}$ depends at least on one of the causal variables in $\mathcal{M}$. If one of the causal variables in $\mathcal{M^{\prime}}$ depended on more than one causal variable in $\mathcal{M}$, it would create dependency between two variables in $\mathcal{M^{\prime}}$ and violate the above equality. Therefore, no variable in $\mathcal{M^{\prime}}$ depends on more than one causal variable in $\mathcal{M}$. Consequently, the transformation $\phi_{\mathcal{Z}}$ is a component-wise transformation.
@@
(Identifiability of latent causal models.) Let $\mathcal{M}=(\mathcal{A},\mathcal{X},g,\mathcal{I})$ and $\mathcal{M}^{\prime}=(\mathcal{A}^{\prime},\mathcal{X},g^{\prime},\mathcal{I})$ be two LCMs with shared observation space $\mathcal{X}$ and shared intervention targets $\mathcal{I}$. Suppose the following conditions are satisfied:
1. Identical correspondence assumptions explained in 3.1.
2. Soft interventions satisfy Assumption 3.3.
3. The causal and exogenous variables are real-valued.
4. The causal and exogenous variables follow a multivariate normal distribution.